# colorlinks, linkcolor=red!50!black, citecolor=blue!50!black, urlcolor=blue!80!black

### v1

@@ Abstract
Deep Learning, Transformers, Explainability, Point Clouds, Self-Supervision
@@ I Introduction
As this approach became more common, more sophisticated techniques of self-supervision were devised [5, 6, 7]. However, few studies have actually delved into what it is that the models actually learn [8, 9, 10] and even fewer have examined point clouds in particular [11, 12]. Coupled with several challenges that arise in the domain shift from language/image to point clouds, we find that this particular area is lacking.
@@
To summarize, our contributions are threefold:
@@
  • We perform comparative evaluations of the baseline Masked AutoEncoding pretraining scheme to its main contender, Contrastive Learning, and discover several key differences in their learned representations, such as a complementary reliance on feature and positional embeddings respectively. (Sec. VI)
@@ II-A Deep Learning on Point Clouds
The domain shift from 2D to 3D came with several extra challenges, due to the lack of grid structure and uneven density of point clouds. Early attempts involved the application of image methods as-is, using multi-view images [15, 16] or voxels [17, 18, 19]. Not long after that, specialized architectures emerged. PointNet [20] pioneered point-wise MLPs and pooling for extracting global features. PointNet++ [21] later followed a multi-scale approach by incorporating neighborhood information. Following this paradigm, other works create more complex kernels, taking advantage of geometric priors [22, 23, 24, 25, 26].
More recently, transformer variants have been successfully applied to the point cloud domain. Originally introduced in [27], transformers were quickly established as the go-to choice in NLP. Dosovitskiy et al.[13], later introduced the vision transformer, operating on image data. This was accomplished by tokenizing the input into patches and incorporating spatial information through a positional embedding. Point Transformer [28] is among the first works to apply a transformer-like architecture to point cloud data. They apply intra-neighborhood attention to create a patch feature vector and use pooling operations to downsample the point cloud.
@@ II-B Self supervised Learning
The idea of making use of large amounts of unlabeled data to improve a learning model is not new, rather it has been around for decades [34, 35, 36, 37]. Its popularity did not surge until the mid-2010s [38, 2, 3] however, when the availability of unlabeled data at scale started to increase and the potency of hardware could finally keep up to the challenge. Of particular importance is masked autoencoding, pioneered by [1], which revolutionized language modeling. This paradigm was later followed by other works, such as [39, 40, 41], that further scaled up in terms of parameter and dataset size, achieving state-of-the-art results in the few-shot and zero-shot settings.
@@
Despite the scarcity of point cloud data compared to image and language data, and the corresponding need for good self-supervision techniques in this domain, the topic has only recently begun attracting attention. [49] proposes to split point clouds into parts and use these parts to parameterize gaussian mixture models. The model is then trained by using a loss that resembles likelihood maximization. In [50] the authors perform patch-level contrastive learning. By rotating a query patch they form positive pairs, while negative pairs are formed by taking into account any other patch of the shape. In [51], a similar approach to masked autoencoding is followed, that is, completion of occluded point clouds acquired by taking custom viewpoints through virtual cameras. [52] argues that performing pretext tasks on single object point clouds might have limited benefits in real world applications. They instead sample positive and negative pairs from complex scene scans, hoping to get a better estimate of the target distributions, and apply a contrastive loss to train their model.
@@ III Baseline Method, Tools, and Dataset
In this section, we briefly overview our baseline model and pretraining setup. After training and finetuning, we assess the quality of the learned features both qualitatively and quantitatively by using explainability tools and measuring the accuracy in the classification task. Further details regarding the model and the training process can be found in the appendix. We utilize pytorch lightning [56] for our codebase, in favor of clarity, transparency, and reproducibility. The code is publicly available on github [^1] https://github.com/VVRPanda/ExplainableTransformerPC.
@@ III-C Dataset
As opposed to their image counterparts, point cloud datasets are generally fewer and significantly smaller in size. Additionally, there are a lot of impactful differences between point cloud datasets, including sampling density, scanning device-specific artifacts, and shape variations in general (real-world objects, CAD models). As a result, a model trained on a specific dataset may not generalize to other data distributions. Our first goal towards explainability is to see how the amount of training data might affect the network’s performance. To this end, we concatenate the commonly used ShapeNet [57] with the more obscure CC3D [58] (41k and 43k training samples, respectively). CC3D contains high resolution, single-object, fine-grained CAD shapes from arbitrary categories. To study the effects of the increased volume of data, we conduct pretraining experiments using only ShapeNet (S) and the concatenated dataset (C+S). We compare the two by finetuning for classification in ModelNet40 and ScanObjectNN and report the results in Tables I, II. In both cases, our proposed dataset helps achieve better accuracy. For the rest of the paper, we will be using C+S for pretraining, unless otherwise specified.
@@ IV Strategic Unfreezing
TABLE I: Comparison of different transformer based models and pretraining schemes without using voting on the ModelNet40 test set.
Model Acc
Point-BERT [31] 92.7
Point-MAE [4] 93.2
PCT-2L [29] 93.2
PCT-3L [29] 93.4
Point-M2AE [59] 93.4
Ours (S) 93.6
Ours (C+S) 93.7
TABLE II: Classification results on ScanObjectNN. The HARD variant refers to PB-T50-RS that is the most challenging one. SoNN indicates that the model has been pretrained for a few epochs in the ScanObjectNN-hard dataset (Bold and underlined denotes 1st and 2nd highest respectively).
@@
Point-BERT [31] 87.43 88.12 83.07
Point-MAE [4] 90.02 88.29 85.18
Ours(S - unfr 200) 90.19 88.98 84.98
@@
Ours (SoNN) 90.36 89.67 85.25
@@ V-B Attention Distance
$mAD_{l,h}=\frac{1}{N^{2}}\sum_{ij}A_{l}\odot D$
Where $A_{l}^{N\times N}$ is the attention matrix of layer $i$ and $D^{N\times N}$ is the distance matrix, where each entry $d_{i,j}$ is the Euclidean distance between centers $i$ and $j$ and the $\odot$ symbol denotes the Hadamard product. The product is averaged across all entries to obtain a mean attention-distance value, which is also averaged across the validation set of ModelNet40.
Based on the graphs in Figure 6, it can be inferred that with the increase in the amount of training data, the earlier layers tend to focus more locally with certain attention heads, while others still maintain a global focus. This observation suggests that the transformer network may begin to learn the inductive bias of the convolution, that is, attending to local features, while still retaining the capability to aggregate global information, thereby resulting in more comprehensive representations.
Interestingly, our findings are consistent with those of [8, 60], indicating that transformers may exhibit akin behaviors across domains. A noticeable difference is that in the case of ShapeNet (Figure 6b) the first two layers seem attend globally with all of their attention heads, whereas when CC3D is added to the pretraining data, this phenomenon is only observed in the first layer (Figure 6c). We assume that the network requires a rough understanding of the entire shape before opting to extract local features. This demeanor should be studied further if point cloud datasets at large scale become available in the future.
@@ VI Contrastive Learning
As opposed to masked autoencoding, an interesting property of contrastive learning is that it allows the classification token to be trained during the pretraining stage. In fact, by examining the attention scores, we notice that the classification token attends to characteristic areas of the input shape, and exhibits symmetry, by attending to shape parts with similar semantics (Figure 8). We also observe that from layers 8 and onward, the CLS token attends uniformly to the whole shape. This is verified by Figure 9b where layers 10 and 11 have the exact same attention distance across all heads. Finally, we see that through MoCo pretraining the backbone does not learn to attend to neighboring patches, indicating that the training scheme might require more data or alterations to converge. For the sake of completeness, we also tested a MoCo and MAE hybrid (details in the supplementary material).
@@
VII Discussion
Using our proposed unfreezing strategy, we manage to outperform our baseline model, as well as models that use transformers in a similar way. We do not use voting for any of the results presented in this paper, nor do we include random transforms in the validation set. The method we suggest combined with our learning rate scheduling policy yield stable training with very close convergence points and weaker fluctuations.
@@ VII Discussion
In our explainability experiments, we observe that point neighborhoods with sharp geometric features are generally more likely to be attended to by the classification token. This is an indicator of good pretraining, as sharp features are more relevant descriptors of the sample’s class. We also notice that the transition from local to global seems characteristic of transformers. This property is experimentally backed in the image domain [8], and our own experiments confirm that this holds true in point clouds as well.
@@
VIII Future Work
Reasoning about the learned representations requires depth and repeating all finetuning experiments for each new pretrained network. In future work, we plan on experimenting and associating the properties of pretrained networks with specific downstream tasks. We theorize that due to their distinct characteristics, some pretraining schemes may offer greater advantages for a specific downstream task than others.

### v3

# This work has received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101092875 - DIDYMOS-XR: Digital DynaMic and responsible twinS for XR.)

@@ Abstract
Deep Learning, Explainability, Point Clouds, Self-Supervision, Transformers
[…]
=-21pt
@@ I Introduction
As this approach became more common, more sophisticated techniques of self-supervision were devised [4, 5, 6]. However, few studies have actually delved into what it is that the models actually learn [7, 8, 9] and even fewer have examined point clouds in particular [10, 11]. Point Cloud models differ significantly from image-based ones. As opposed to an image’s canonical grid structure, uniform density and general data availability, the point clouds are highly irregular, highly inconsistent in terms of density, and are generally scarce compared to images. Additionally, contrary to an image’s fixed pixel position, point clouds are unordered, often different in terms of cardinality, and require that their handling be invariant to permutations. Due to these challenges that arise in the domain shift from language/image to point clouds, we find that this particular area is lacking.
@@
To summarize, our contributions are twofold:
@@ II-A Deep Learning on Point Clouds
The domain shift from 2D to 3D came with several extra challenges, due to the lack of grid structure and uneven density of point clouds. Early attempts involved the application of image methods as-is, using multi-view images [14, 15] or voxels [16, 17]. Not long after that, specialized architectures emerged. PointNet [18] pioneered point-wise MLPs and pooling for extracting global features. PointNet++ [19] later followed a multi-scale approach by incorporating neighborhood information. Following this paradigm, other works create more complex kernels, taking advantage of geometric priors [20, 21, 22, 23, 24].
More recently, transformer variants have been successfully applied to the point cloud domain. Originally introduced in [26], transformers were quickly established as the go-to choice in NLP. Dosovitskiy et al.[12], later introduced the vision transformer, operating on image data. This was accomplished by tokenizing the input into patches and incorporating spatial information through a positional embedding. This approach also allowed data from different modalities to be used jointly [27, 28], making the transformer the most prominent architecture in that field. Point Transformer [29] is among the first works to apply a transformer-like architecture to point cloud data. They apply intra-neighborhood attention to create a patch feature vector and use pooling operations to downsample the point cloud.
@@ II-B Self supervised Learning
The idea of making use of large amounts of unlabeled data to improve a learning model is not new, rather it has been around for decades [35, 36, 37, 38]. Its popularity did not surge until the mid-2010s [39, 2, 3] however, when the availability of unlabeled data at scale started to increase and the potency of hardware could finally keep up to the challenge. Of particular importance is masked autoencoding, pioneered by [1], which revolutionized language modeling. This paradigm was later followed by other works, such as [40, 41, 42], that further scaled up in terms of parameter and dataset size, achieving state-of-the-art results in the few-shot and zero-shot settings. In [43] they experiment with pretraining by using synthetic data, artificially injected with desirable properties.
@@
Despite the scarcity of point cloud data compared to image and language data, and the corresponding need for good self-supervision techniques in this domain, the topic has only recently begun attracting attention. [51] proposes to split point clouds into parts and use these parts to parameterize gaussian mixture models. The model is then trained by using a loss that resembles likelihood maximization. In [52] the authors perform patch-level contrastive learning. By rotating a query patch they form positive pairs, while negative pairs are formed by taking into account any other patch of the shape. A similar approach is followed in [53], finding positive and negative pairs by using an inductive model to perform pseudolabeling. In [54], a similar approach to masked autoencoding is followed, that is, completion of occluded point clouds acquired by taking custom viewpoints through virtual cameras. [55] argues that performing pretext tasks on single object point clouds might have limited benefits in real world applications. They instead sample positive and negative pairs from complex scene scans, hoping to get a better estimate of the target distributions, and apply a contrastive loss to train their model.
@@ III Baseline Method, Tools, and Dataset
In this section, we briefly overview our baseline model and pretraining setup. After training and finetuning, we assess the quality of the learned features both qualitatively and quantitatively by using explainability tools and measuring the accuracy in the classification task. Further details regarding the model and the training process can be found in the appendix. We utilize pytorch lightning [^1] https://www.pytorchlightning.ai/ for our codebase, in favor of clarity, transparency, and reproducibility. The code is publicly available on github [^2] https://github.com/VVRPanda/ExpPoint-MAE.git.
@@ III-C Dataset
As opposed to their image counterparts, point cloud datasets are generally fewer and significantly smaller in size. Additionally, there are a lot of impactful differences between point cloud datasets, including sampling density, scanning device-specific artifacts, and shape variations in general (real-world objects, CAD models). As a result, a model trained on a specific dataset may not generalize to other data distributions. Our first goal towards explainability is to see how the amount of training data might affect the network’s performance. To this end, we concatenate the commonly used ShapeNet [59] with the more obscure CC3D [60] (41k and 43k training samples, respectively). CC3D contains high resolution, single-object, fine-grained CAD shapes from arbitrary categories. To study the effects of the increased volume of data, we conduct pretraining experiments using only ShapeNet (S) and the concatenated dataset (C+S). We compare the two by finetuning for classification in ModelNet40 [61] and ScanObjectNN [62] and report the results in Tables I, II. In both cases, our proposed dataset helps achieve better accuracy. For the rest of the paper, we will be using C+S for pretraining, unless otherwise specified.
@@ IV Strategic Unfreezing
TABLE I: Comparison of different transformer based models and pretraining schemes with and without using voting on the ModelNet40 test set.
MODEL W/O VOTING VOTING
Point-BERT [32] 92.7 93.2
MaskPoint [63] - 93.8
Point-MAE [25] 93.2 93.8
PCT-2L [30] 93.2 -
PCT-3L [30] 93.4 -
Point-M2AE [64] 93.4 94.0
Ours (S) 93.6 -
Ours (C+S) 93.7 94.2
Blue and Cyan denote 1st and 2nd highest.
TABLE II: Classification results on ScanObjectNN. The HARD variant refers to PB-T50-RS that is the most challenging one.
@@
Point-BERT 87.43 88.12 83.07
MaskPoint 89.3 89.7 84.6
Point-MAE 90.02 88.29 85.18
Point-M2AE 91.22 88.81 86.43
@@
Ours (SoNN11footnotemark: 1) 90.36 89.67 85.25
Blue and Cyan denote 1st and 2nd highest.
11footnotemark: 1: SoNN indicates that the model has been pretrained for a few epochs in the ScanObjectNN-hard dataset.
@@ V-B Attention Distance
$mAD_{l,h}=\frac{1}{N^{2}}\sum_{ij}\mathbf{A_{l}}\odot\mathbf{D}$
Where $\mathbf{A_{l}}^{N\times N}$ is the attention matrix of layer $i$ and $\mathbf{D}^{N\times N}$ is the distance matrix, where each entry $d_{i,j}$ is the Euclidean distance between centers $i$ and $j$ and the $\odot$ symbol denotes the Hadamard product. The product is averaged across all entries to obtain a mean attention-distance value, which is also averaged across the validation set of ModelNet40.
Based on the graphs in Figure 6, it can be inferred that with the increase in the amount of training data, the earlier layers tend to focus more locally with certain attention heads, while others still maintain a global focus. This observation suggests that the transformer network may begin to learn the inductive bias of the convolution, that is, attending to local features, while still retaining the capability to aggregate global information, thereby resulting in more comprehensive representations. This mitigates the general struggle of self-attention to model local relationships [65], without requiring any modifications to the architecture.
Interestingly, our findings are consistent with those of [7, 66, 67], indicating that transformers may exhibit akin behaviors across domains. A noticeable difference is that in the case of ShapeNet (Figure 6b) the first two layers seem attend globally with all of their attention heads, whereas when CC3D is added to the pretraining data, this phenomenon is only observed in the first layer (Figure 6c). We assume that the network requires a rough understanding of the entire shape before opting to extract local features. This demeanor should be studied further if point cloud datasets at large scale become available in the future.
@@ VI Contrastive Learning
As opposed to masked autoencoding, an interesting property of contrastive learning is that it allows the classification token to be trained during the pretraining stage. In fact, by examining the attention scores, we notice that the classification token attends to characteristic areas of the input shape, and exhibits symmetry, by attending to shape parts with similar semantics (Figure 8). We also observe that from layers 8 and onward, the CLS token attends uniformly to the whole shape. This is verified by Figure 9b where layers 10 and 11 have the exact same attention distance across all heads. Finally, we see that through MoCo pretraining the backbone does not learn to attend to neighboring patches, indicating that the training scheme might require more data or alterations to converge. For the sake of completeness, we also tested a MoCo and MAE hybrid (details in the Appendix).
@@ VII Ablation Studies
VIII Discussion
Using our proposed unfreezing strategy, we manage to outperform our baseline model, as well as models that use transformers in a similar way. We do not use voting unless explicitly mentioned for fair comparisons, nor do we include random transforms in the validation set. The method we suggest combined with our learning rate scheduling policy yield stable training with very close convergence points and weaker fluctuations.
@@ VIII Discussion
In our explainability experiments, we observe that point neighborhoods with sharp geometric features are generally more likely to be attended to by the classification token. This is an indicator of good pretraining, as sharp features are more relevant descriptors of the sample’s class. We also notice that the transition from local to global seems characteristic of transformers. This property is experimentally backed in the image domain [7], and our own experiments confirm that this holds true in point clouds as well. An additional surprising finding is the contrast between the behaviors of two finetuned models, one that has been pretrained and one that has not. In the first case the human eye can clearly capture a transition from local to global, whereas in the second case the attention scores are mostly random, containing no identifiable pattern.
@@
IX Future Work
Our analysis in section VI demonstrates that contrastive learning exhibits useful properties, namely, symmetry and the ability to train a classification token in the pretraining stage. Consequently, we believe that contrastive learning can potentially outperform MAE in several tasks, given enough training time and data. In fact, we theorize that due to their distinct characteristics, some pretraining schemes may offer greater advantages for a specific downstream task than others. In future work, we plan on experimenting on a framework for determining which pretraining scheme to use, in order to utilize our models to their fullest potential.
Appendix A MoCo & MAE Hybrid Pretraining
We devise a pretraing sceme that combines both core concepts from MoCo and MAE. The main idea is that the reconstruction loss from MAE will help the backbone converge faster, being a more straightforward task, and the contrastive loss will help in training the CLS token and slightly contribute to the backbone features as well. The scheme is realized through a weighted sum of both losses, appropriately scaled so that they are in the same order of magnitude.
Specifically, the standard reconstruction loss presented in [33] is applied ($L_{rec}$) and a contrastive loss, $L_{con}$, as in [20] is added, using a queue size of $4096$ encoded samples. The final loss function is realized as:
$L=L_{rec}+w_{c}\cdot L_{con}$
Where $w_{c}$ was set to $10^{-2}$ to bring the two losses to the same order of magnitude. We also experimented with a smaller weight($10^{-3}$), as the MoCo loss is responsible mainly for the CLS token. Both approaches, however, produced similar results.
Although the hybrid model’s accuracy reaches an unimpressive $91.8$% score, we believe that with further research it is possible to create a model that makes the best of both worlds (masked autoencoding and contrastive learning).
Figure 13: Confusion matrix visualization for the validation set of ModelNet40. Each row has been normalized through division by the number of samples of that class.
A-A Strategic unfreezing
As briefly discussed in the main paper (due to space constraints), the finetuning data from ModelNet40 have already been seen during the pretraining stage and therefore the distribution of said data has been captured by the backbone. Therefore, it is sufficient to only train the backbone for the last 50 epochs, where the learning rate is small enough so that no overblown weight updates happen.
The ScanObjectNN is probably an even better indicator of this strategy. We observe that in table II, when training on objects without background the same unfreezing point yields the best results. On the other hand, when background is included, unfreezing at 200 epochs is optimal and obtains even better accuracy. While the higher accuracy can easily be attributed to the extra information contained in the background itself, it is important to notice that the background changes the data distribution significantly. This change, coupled with the fact that the data has been obtained through scanning, leads us to think that 50 epochs is no longer enough for the backbone to adjust to the new distribution.
This argument is further verified by finetuning on the hard variant of ScanObjectNN. Sure enough, 100 epochs of training with the unfrozen backbone makes a staggering $+1.18\%$ difference in accuracy compared to 50 epochs, however it is still far from the optimal, which is to pretrain the network for a few epochs on this dataset. In our experiments, 40 epochs of pretraining are adequate to reach an accuracy of $85.25\%$, that surpasses our baseline Point-MAE.
Generally speaking, there is no explicit way to determine the exact epoch, at which the unfreezing yields optimal results. However, our proposed ”strategic unfreezing” method is a practical way of iteratively improving the model at hand, by providing concrete proof regarding the similarity with the original pretrained version of the model. Consequently, it allows the researcher to make informed design choices and gain an overall deeper understanding.
A-B Optimization and learning rate policy
For both our optimizer and learning rate we follow one of [33]’s suggestions. For our finetuning we use the AdamW optimizer with a peak learning rate of $\lambda=5\cdot 10^{-4}$ and a weight decay of $W_{d}=0.05$. We utilize the cosine annealing learning rate scheduler, updating during every step for a total training of 300 epochs. A linear warmup is also performed, with an initial learning rate $\lambda_{i}=10^{-6}$ for a duration of 10 epochs in total. The exact same setup is also used for pretraining, with the exception of peak learning rate, being $\lambda=10^{-3}$. The batch size is 128 for pretraining and 32 for finetuning.
A-C Data Preprocessing
All data samples are normalized to be inside the unit sphere. For data augmentations we apply random sampling of the point clouds and anisotropic scale, following the paradigm of our baseline.
A-D Hardware
All models were trained on a single NVIDIA RTX 4090 GPU. It should be noted that the memory requirements did not exceed 16GB. The training times were (approximately) 16 hours for pretraining on the combined dataset. For finetuning on our best model (unfreeze 250) the training time was 2 hours on ModelNet40 and 1 hour on ScanObjectNN. Note that when unfreezing at an earlier stage, the total time is lengthened due to the weight updates happening inside the backbone. It takes the model approximately $36$ ms to process a batch with 16 samples.
Appendix B Class Clusters
Figure 14: Visualization of samples from problematic classes of ModelNet40. First, second and third rows correspond to ”flower pot”, ”plant” and ”vase” respectively.
In order to get a better idea of the causes behind misclassified objects we present a confusion matrix (figure 13) generated from samples of ModelNet40 validation set. The most striking entry is the 15th category, ”flower pot”, being mistaken for classes 26, ”plant”, and 37, ”vase”. These categories contain an unreasonably large amount of overlap, e.g. pots have plants, plants have pots and a lot of pots have a vase-like shape (figure 14). Therefore, it is hardly a surprising result that these objects are being misclassified, and this particular error is not a good descriptor of the model’s performance.
Figure 15: Clustering performed by the model in the feature space for 10 arbitrary classes. The model has been pretrained using the MAE pipeline. T-SNE is used to project the models two the 2-dimensional space. (A) A sample from the airplane class is placed near the guitar cluster. An ”excusable” choice as both share a lengthy shape and sharp protruding parts. (B) Two shape from the person and plant classes are placed further from their clusters, and near each other, as they have a near non-existent depth. (C) Two uniquely shaped lamps have been assigned to a cluster of their own.
Figure 15: Clustering performed by the model in the feature space for 10 arbitrary classes. The model has been pretrained using the MAE pipeline. T-SNE is used to project the models two the 2-dimensional space. (A) A sample from the airplane class is placed near the guitar cluster. An ”excusable” choice as both share a lengthy shape and sharp protruding parts. (B) Two shape from the person and plant classes are placed further from their clusters, and near each other, as they have a near non-existent depth. (C) Two uniquely shaped lamps have been assigned to a cluster of their own.
Figure 16: Clustering performed by the model in the feature space for 10 arbitrary classes. The model has been pretrained using the MoCo pipeline. T-SNE is used to project the models two the 2-dimensional space. (A) A tall cactus is clustered together with the people class. (B1), (B2), Two lamps with distinct characteristics have been assigned to a separate cluster. (C) A plant with hanging leaves is placed along with the table cluster, most likely because the leaves share a similar structure with table legs.
In figures 16 and 16 (MAE and MoCo respectively) we visualize the feature vectors of 15 shapes from 10 arbitrary classes randomly sampled from the validation set of ModelNet40 and projected onto the 2d space by using T-SNE. Naturally, the test accuracy is the sole indicator of how well the clustering is performed. However, by using this technique we are hoping to understand whether the clustering/classification operates within the realms of human logic. For this purpose, we are mostly interested in the misclustered samples.
All in all, these mistakes seem plausible from a human perspective. In particular, the model seems to be ”fooled” by objects that have similar height/width or are composed by roughly equivalent parts. This is hardly a fault however, as it only decreases the score on the ModelNet40 benchmark, whilst achieving greater generalization ability.